# Adaptive Rank, Reduced Forgetting: Knowledge Retention in Continual Learning Vision-Language Models with Dynamic Rank-Selective LoRA

### v5

@@ 3.2 Analyses on LoRA Location and Rank
Figure 3: Evaluations of downstream task adaptation and zero-shot capability retention when training CLIP with varying LoRA insertion points and ranks. Shapes indicate encoders, colors denote transformer modules, and sizes reflect rank values.
Figure 4: Analysis of the impact of removing trained LoRA modules on downstream task adaptation (Aircraft) and retention of pre-trained capabilities (ImageNet-1k).
@@
To achieve efficient learning, we explore LoRA and aim to optimize its usage strategy. To analyze the impact of LoRA’s rank and placement on downstream learning and forgetting, we conduct a series of experiments: (1) We apply LoRA components with varying ranks at different locations in the pre-trained CLIP model (Fig. 4); (2) We apply LoRA across all PTM weights and systematically prune them for evaluation (Fig. 4). We train the model on downstream datasets and assess its performance on both (1) the new task and (2) an additional reference dataset to evaluate the preservation of pre-trained capabilities.
@@
Applying LoRA on all weights. As shown in Fig. 4 and Fig. 4, applying LoRA to specific modules alone may limit learning capacity in different ways: (1) The different LoRA placement locations significantly affect downstream performance and the forgetting/retention of pre-trained capabilities; (2) Updating only the vision encoder leads to more forgetting of pre-trained capabilities; (3) Updating only the attention modules helps mitigate forgetting but limits the learning ability on new tasks. These observations align with the distinct roles of different PTM components. Rather than manually selecting LoRA placement, we apply it across all weights and optimize for an adaptive configuration.
@@
The effects of different LoRA ranks on balancing learning and forgetting. From Fig. 4, we can observe that, despite learning low-rank space, a relatively higher-rank LoRA facilitates learning new tasks but tends to increase forgetting, while lower-rank LoRA mitigates forgetting but limits adaptation. This suggests that a lower-rank LoRA produces stronger regularization in maintaining the learned knowledge by sacrificing the learning of new knowledge. Importantly, we identify a balance at a moderately small rank, which reduces forgetting while maintaining adaptability. This implies that minimizing the rank while fitting on new data may find an optimized trade-off for CL.
The balance varies by location and task. The ranks for a good balance vary by placement (Fig. 4), and pruning LoRA at different locations has varying effects (Fig. 4), indicating that not all LoRA components contributed equally for a task. For example, removing updates in the Value projection of the Attention module in the vision encoder minimally affected adaptation but significantly restored pre-trained performance. Interestingly, pruning the MLP layer of the vision encoder slightly reduced task performance but enhanced pre-trained capabilities. These findings motivate our adaptive approach to dynamically determine ranks based on location and task.
3.3 Proposed CoDyRA
@@
Overview. Our analyses in Sec. 3.2 suggest applying LoRA across all CLIP modules while adaptively selecting an appropriate rank for each task and location to balance learning and forgetting. While finding an ideally global optimal rank is challenging, we approach it by jointly optimizing downstream task performance and minimizing the number of ranks using sparsity-promoting regularization.
@@
We introduce dynamic rank-selective parameter updates for all pre-trained weight matrices within the Attention and MLP modules of both the vision and text encoders, as shown in Fig. 2. These updates are adaptively inserted based on the importance of each rank for the current training data. For clarity, during training on task $t$, we denote the pre-trained weights—or those updated from the previous tasks—as $\{\mathbf{W}^{t,m}_{0}\}_{m=1}^{M}$, where $M$ represents the total number of pre-trained weight matrices augmented with LoRA. After training on each task $t$, these dynamic rank updates are merged back into the original pre-trained weights.
@@
Low-rank updates with rank importance. For efficient adaptation, instead of directly updating a pre-trained weight matrix $\mathbf{W}_{0}^{t,m}\in\mathbb{R}^{d\times k}$, LoRA [21] introduces two low-rank matrices, $\mathbf{B}^{t,m}\in\mathbb{R}^{d\times r}$ and $\mathbf{A}^{t,m}\in\mathbb{R}^{r\times k}$, such that $\Delta\mathbf{W}=\mathbf{B}^{t,m}\mathbf{A}^{t,m}$, where $r\ll\min(d,k)$. The updated weight matrix is then defined as:
@@
In this formulation, the original weights $\mathbf{W}_{0}^{t,m}$ remain fixed, and only $\mathbf{B}^{t,m}$ and $\mathbf{A}^{t,m}$ are trained, reducing the number of trainable parameters to $r(d+k)$. Notably, LoRA incurs no additional computational overhead during inference, as the low-rank updates can be efficiently merged into the pre-trained weights, as shown in Eq. (1).
@@
To automatically and adaptively decide the rank in the LoRA updates, we introduce a learnable importance weight vector $\mathbf{w}^{t,m}\in\mathbb{R}^{r}$ and incorporate it into the formulation of LoRA as follows:
$\Delta\mathbf{W}^{t,m}=\sum\nolimits_{i=1}^{r}\mathbf{w}_{i}^{t,m}\mathbf{B}_{i}^{t,m}\mathbf{A}_{i}^{t,m},$ (2)
where $\mathbf{w}^{t,m}$ is learnable and represents the importance of each rank. The formulation is based on a reformulation of LoRA updates in Eq. (1) based on Singular Value Decomposition (SVD) [64, 12, 32, 38, 63]. The importance weights are corresponding to the diagonal elements in the singular value matrix of SVD.
Introducing importance weights incurs minimal overhead, adding only $r$ additional parameters per module. For initialization, we follow the standard convention: matrix $\mathbf{A}$ is randomly initialized, matrix $\mathbf{B}$ is set to zero, and the importance weights $\mathbf{w}$ are randomly initialized.
@@
Table 1: Comparisons on X-TAIL for each domain in terms of “Transfer”, “Average”, and “Last” scores (%). The best and the second best results are highlighted in red and blue, respectively. Methods marked with $\dagger$ indicate the use of domain prediction or distribution detection [61, 58].
@@
    Method Aircraft [35] Caltech101 [15] DTD [8] EuroSAT [19] Flowers [40] Food [5] MNIST [11] OxfordPet [41] Cars [26] SUN397 [57] Average
@@
   Zero-shot [42] 76.8 37.3 36.7 63.6 84.0 46.7 86.7 66.1 63.7 62.4
   LwF [29] 66.6 26.9 19.5 51.0 78.4 26.6 68.9 35.5 56.1 47.7
   WiSE-FT [54] 70.1 31.9 25.3 56.3 79.8 29.9 74.9 45.6 56.8 52.3
   iCaRL [43] 71.7 35.0 43.0 63.4 86.9 43.9 87.8 63.7 60.0 61.7
   ZSCL [66] 73.3 32.6 36.8 62.1 83.8 42.1 83.6 56.5 60.2 59.0
@@
   LwF [29] 24.7 79.7 38.3 36.9 63.9 81.0 36.5 71.9 42.7 56.7 53.2
   WiSE-FT [54] 27.1 76.5 40.9 31.3 68.7 81.6 31.4 74.7 51.7 58.4 54.2
   iCaRL [43] 25.4 72.1 37.5 51.6 65.1 87.1 59.1 88.0 63.7 60.1 61.0
   ZSCL [66] 36.0 75.0 40.7 40.5 71.0 85.3 46.3 83.3 60.7 61.5 60.0
@@
   LwF [29] 25.5 72.1 38.9 55.4 65.5 87.3 81.9 88.6 63.6 61.5 64.0
   WiSE-FT [54] 21.8 76.8 42.9 20.8 77.5 84.9 30.7 76.6 75.8 72.5 58.0
   iCaRL [43] 25.5 72.1 38.9 55.4 65.5 87.3 81.9 88.6 63.6 61.5 64.0
   ZSCL [66] 33.1 75.3 43.5 35.2 74.6 87.4 50.4 84.2 77.3 73.4 63.4
@@
Dynamic rank-selective updates with sparse rank importance. The formulation in Eq. (2) enables the model to dynamically adjust the importance of each rank through gradient descent. [^1] To align the training process with the pre-trained model, we use CLIP’s contrastive loss as the optimization objective during pre-training. However, not all ranks contribute equally to downstream task performance (as shown in Fig. 4). To ensure that only the most impactful ranks are retained for learning on new task while removing unnecessary ranks to mitigate forgetting of already-learned knowledge, we propose optimizing the importance weights with a sparsity-promoting regularization through minimizing the $\ell_{1}$ norm. The optimization objective for parameters $\{\mathbf{w}^{t,m},\mathbf{B}^{t,m},\mathbf{A}^{t,m}\}_{m=1}^{M}$ learned on task $t$ is:
$\mathcal{L}_{\text{train}}^{t}:=\mathcal{L}_{\text{sup}}^{t}+\lambda\sum\nolimits_{m=1}^{M}\|\mathbf{w}^{t,m}\|_{1},$ (3)
where $\mathcal{L}_{\text{sup}}^{t}$ is the supervised training loss, $M$ is the total number of pre-trained weight matrices inserted $r$-rank update, and $\mathbf{w}^{t,m}$ represents the importance weights of each rank added to the weight matrix $\mathbf{W}^{t,m}_{0}$. The $\ell_{1}$ regularization strength is controlled by $\lambda$.
To handle the non-differentiable $\ell_{1}$ regularization applied to the importance weights, we adopt the proximal gradient method [4]. The $i$-th element of $\hat{\mathbf{w}}^{t,m}$ for rank $i$ is updated via soft-thresholding:
where $\hat{\mathbf{w}}^{t,m}_{i}$ denotes the value of ${\mathbf{w}^{t,m}_{i}}$ after applying the gradient update from the supervised loss $\mathcal{L}_{\text{sup}}^{t}$. The threshold parameter $\kappa$ is initially set to zero and gradually increases to a maximum value $\kappa_{\text{max}}$, analogous to linearly scheduling the regularization strength $\lambda$. The indicator function $\mathbbm{1}(\cdot)$ returns 1 if the condition is met and 0 otherwise, while $\mathrm{sign}(\cdot)$ represents the +/+/-+ / - sign of the input.
This approach preserves only the ranks with significant importance, pruning those with low relevance. By iteratively adjusting weightings, the model focuses dynamic parameter updates on the most relevant components. During the early stages of training, dense updates are performed without Eq. (4) to allow all ranks to capture task-relevant information before applying sparsity constraints.
Continually enhancing the pre-trained model. After training, ranks with zero importance weights are pruned, while non-zero ranks are merged into the pre-trained weights by substituting Eq. (2) into the formulation Eq. (1). When a new task is introduced, a fresh set of CoDyRA modules is initialized, and computations involve only these newly initialized CoDyRA modules alongside the continually updated pre-trained model. This approach enables the model to efficiently leverage accumulated knowledge while maintaining a minimal computational footprint, ensuring no additional overhead.
Table 2: Comparisons on 5-shot MTIL setting, evaluated in a similar manner as in Table 1.
    Method Aircraft [35] Caltech101 [15] CIFAR100 [27] DTD [8] EuroSAT [19] Flowers [40] Food [5] MNIST [11] OxfordPet [41] Cars [26] SUN397 [57] Average

### v6

@@ 3.2 Analyses on LoRA Location and Rank
To enable efficient and natural continual updating of PTMs, we employ LoRA and analyze its learning and forgetting capability for optimizing its usage. We examine how LoRA’s rank and placement affect downstream learning and forgetting through two types of analyses (Pearl, 2022; Meng et al., 2022), providing a comprehensive view of their overall impact. (1) We first apply LoRA components with varying ranks at different locations in the pre-trained CLIP, for measurement of importance or contribution (Fig. 4); (2) We apply LoRA across all PTM weights and systematically prune them for evaluation, for testing the standalone capability (Fig. 4). We train the model on downstream datasets and assess its performance on both (1) the new task, i.e., learning new knowledge, and (2) an additional reference dataset to evaluate the retention and forgetting of existing capabilities. We achieve three main results to guide methodology design. Without loss of generality, the learning–forgetting analyses can be extended to each learning step in straightforward, assumption-free CL.
@@
Applying LoRA on all weights. Unlike prior works (Hu et al., 2021; Zhang et al., 2024a; Meng et al., 2024) that only apply LoRA to limited components by default, we first examine the impact of different LoRA placements. As shown in Fig. 4 and Fig. 4, applying LoRA to only specific modules alone may limit learning capacity in different ways: (1) LoRA placement strongly impacts downstream performance and retention of pre-trained capabilities; (2) Updating only the vision encoder causes more forgetting; (3) Updating only attention modules mitigates forgetting but limits learning on new tasks. These observations reflect the distinct roles of PTM components. Takeaway 1: Instead of manual or fixed LoRA placement at specific parameters of PTM like previous methods (Hu et al., 2021; Liang & Li, 2024; Yu et al., 2024; Zhang et al., 2024a; Meng et al., 2024), we apply LoRA to all weights and optimize for an adaptive configuration.
@@
The effects of LoRA ranks on balancing learning and forgetting. We identified that the rank number of LoRA plays a critical role in balancing learning and forgetting. As shown in Fig. 4, we can observe that a relatively higher-rank LoRA facilitates learning new tasks but tends to increase forgetting, while a lower-rank LoRA mitigates forgetting but limits adaptation. This suggests that lower-rank LoRA interferes less with existing knowledge (i.e., reduces forgetting) while learning less new knowledge. Fortunately, we identify a balance (top-right in Fig. 4), where a moderately small rank reduces forgetting while enabling enough adaptability.
During training, low-rank parameter updates tend to enable the updated model to be closer to the original model, with less interference and forgetting. In the extreme case, a rank-zero LoRA learns nothing and introduces no forgetting. Thus, minimizing LoRA rank for each new task in CL can reduce forgetting without extra assumptions. On the other hand, the task-learning objective emphasizes new knowledge acquisition, enabling a balance. Takeaway 2: A plasticity–stability balance exists and associates with LoRA rank (high vs. low), which can be adaptively achieved by jointly optimizing the task objective and minimizing LoRA ranks.
The balance varies by location and task. Fig. 4 shows that the optimal rank for balance depends on both parameter placement and the learning task, consistent with the pruning-based sufficiency analysis in Fig. 4. For example, removing updates to the Value projection in the Attention module of the vision encoder had little impact on adaptation but substantially restored pre-trained performance. Interestingly, pruning the MLP layer slightly reduced task performance while enhancing pre-trained capabilities. Takeaway 3: The learning–forgetting balance point tied to rank varies across modules and tasks, necessitating adaptive optimization.
3.3 Proposed CoDyRA with Sparsity-Induced Rank Selection and Minimization
@@
Overview. Our analyses in Sec. 3.2 suggest applying LoRA across all CLIP modules while adaptively optimizing for an appropriate rank for each specific task and location to balance learning and forgetting. Together with the task learning objective, minimizing active LoRA ranks acts as a regularizer, keeping the updated model closer to its previous state and alleviating forgetting. However, the balanced rank between learning and forgetting is not consistent across parameters or tasks (Takeaway 3). We address this with a sparsity-promoting regularization for the LoRA adapters.
We introduce adaptive rank-selective updates for all pre-trained weight matrices in the Attention and MLP modules of both vision and text encoders (Fig. 2). The sparsity-promoting regularization enables adaptive rank selection and minimization for each LoRA adapter on each task. For each task in CL, beyond adaptive rank minimization, CoDyRA requires no additional priors or assumptions and does not store or maintain information (e.g., data, parameters, or statistics) from previous steps, unlike prior methods (Zheng et al., 2023; Yu et al., 2024; Xu et al., 2024).
@@
During training on task $t$, we denote the pre-trained weights, or those updated from previous tasks, as $\{\mathbf{W}^{t,m}_{0}\}_{m=1}^{M}$, where $M$ is the total number of weight matrices updated by LoRA. After each task, the dynamic rank updates are merged back into the original weights.
@@
Regularized Low-rank updates with rank importance. To update weight matrix $\mathbf{W}_{0}^{t,m}\in\mathbb{R}^{d\times k}$, LoRA (Hu et al., 2021) introduces two low-rank matrices, $\mathbf{B}^{t,m}\in\mathbb{R}^{d\times r}$ and $\mathbf{A}^{t,m}\in\mathbb{R}^{r\times k}$, such that $\Delta\mathbf{W}^{t,m}=\mathbf{B}^{t,m}\mathbf{A}^{t,m}$, where $r\ll\min(d,k)$. The updated weight matrix is then defined as:
@@
Only $\mathbf{B}^{t,m}$ and $\mathbf{A}^{t,m}$ are trained for each task and then merged into the weights. LoRA incurs no additional computational overhead during inference.
As shown in Sec. 3.2, minimizing the active rank for each LoRA adapter alleviates forgetting, which can be formulated as minimizing a regularizer $R(\Delta\mathbf{W}^{t,m})\doteq R(\mathbf{B}^{t,m}\mathbf{A}^{t,m})$. The minimum of $R(\Delta\mathbf{W}^{t,m})$ is zero, i.e., $\min R(\Delta\mathbf{W}^{t,m})=0$, corresponding to a rank-zero matrix—no update to $\mathbf{W}^{t,m}$ and hence no forgetting. By jointly optimizing the new-task learning objective and the minimized $R(\Delta\mathbf{W}^{t,m})$, we achieve a balance between learning and forgetting.
@@
To automatically and adaptively minimize the active ranks in LoRA, we formulate $R(\Delta\mathbf{W}^{t,m})$ through a rank-selection LoRA. We introduce a learnable importance weight vector $\mathbf{w}^{t,m}\in\mathbb{R}^{r}$ to indicate the active ranks for a LoRA adapter:
$\Delta\mathbf{W}^{t,m}=\sum\nolimits_{i=1}^{r}\mathbf{w}_{i}^{t,m}\mathbf{B}_{:,i}^{t,m}\mathbf{A}_{i,:}^{t,m}$ (2)
where $\mathbf{w}_{i}^{t,m}$ is the learnable importance weight associated with rank $i$, $\mathbf{B}_{:,i}^{t,m}$ denotes the $i$-th column of $\mathbf{B}^{t,m}$, and $\mathbf{A}_{i,:}^{t,m}$ denotes the $i$-th row of $\mathbf{A}^{t,m}$. Minimizing $R(\Delta\mathbf{W}^{t,m})$ can be achieved through optimizing $\mathbf{w}^{t,m}$.
Adaptive rank-selective updates with sparse-prompting regularization. The formulation in Eq. (2) enables the model to dynamically optimize the importance of each rank and minimize active ranks through gradient descent. [^1] To align the training process with the pre-trained model, we use CLIP’s contrastive loss as the optimization objective during pre-training. To only retain the essential ranks for learning new tasks (thus reducing forgetting), we can achieve optimizing $R(\Delta\mathbf{W}^{t,m})$ through an $\ell_{1}$ norm-based sparsity-promoting regularization on $\mathbf{w}^{t,m}$, i.e., $\|\mathbf{w}^{t,m}\|_{1}$. The optimization objective for parameters $\{\mathbf{w}^{t,m},\mathbf{B}^{t,m},\mathbf{A}^{t,m}\}_{m=1}^{M}$ learned on task $t$ is:
$\mathcal{L}_{\text{train}}^{t}:={{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathcal{L}_{\text{sup}}^{t}}}+{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\lambda\sum\nolimits_{m=1}^{M}\|\mathbf{w}^{t,m}\|_{1}}},$ (3)
where $\mathcal{L}_{\text{sup}}^{t}$ is the supervised training loss for current task, $M$ is the number of weight matrices inserted $r$-rank update, and $\mathbf{w}^{t,m}$ represents the importance weights of each rank added to the weight matrix $\mathbf{W}^{t,m}_{0}$. The $\ell_{1}$ regularization strength is controlled by $\lambda$.
To handle the non-differentiable $\ell_{1}$ regularization applied to the importance weights, we adopt the proximal gradient method (Beck & Teboulle, 2009). The $i$-th element of $\hat{\mathbf{w}}^{t,m}$ for rank $i$ is updated via soft-thresholding:
where $\hat{\mathbf{w}}^{t,m}_{i}$ denotes the value of $\mathbf{w}^{t,m}_{i}$} after applying the gradient update from the supervised loss $\mathcal{L}_{\text{sup}}^{t}$. The threshold $\kappa$ increases from zero to $\kappa_{\text{max}}$, analogous to linearly scheduling $\lambda$, and adaptively induces sparsity in $\mathbf{w}$ based on the task learning objective. The indicator $\mathbbm{1}(\cdot)$ returns 1 if the condition holds and 0 otherwise, while $\mathrm{sign}(\cdot)$ gives the input’s sign ($\pm$). More details are left in Appendix A.
This approach preserves only the ranks with significant importance for learning the current task and prunes those with low relevance, adapting across weights and tasks. The importance weights on preserved ranks softly highlight their relative significance. Early in training, dense updates are applied without Eq. (4), allowing all ranks to capture task-relevant information before sparsity is enforced.
Continually learn on new tasks. CoDyRA views CL as a sequence of tasks with LoRA, without storing past data Zheng et al. (2023), task information Yu et al. (2024); Xu et al. (2024), or relying on assumptions (Liang & Li, 2024). For each new task, new CoDyRA modules are initialized and optimized; low-importance ranks are pruned, and non-zero ranks merged into the weights. Sparsity-promoting regularization controls task capacity, keeping the model close to its previous state and reducing forgetting. This enables efficient continual updating with only sparsity-based regularization.
@@
Table 1: Comparisons on X-TAIL for each domain. The best and the second best results are highlighted in red and blue, respectively. Methods marked with $\dagger$ indicate the use of domain prediction or distribution detection (Yu et al., 2024; Xu et al., 2024).
Method
@@
Aircraft Caltech101 DTD EuroSAT Flowers Food MNIST OxfordPet Cars SUN397 Average
@@
   Zero-shot (Radford et al., 2021) 76.8 37.3 36.7 63.6 84.0 46.7 86.7 66.1 63.7 62.4
   LwF (Li & Hoiem, 2017) 66.6 26.9 19.5 51.0 78.4 26.6 68.9 35.5 56.1 47.7
   WiSE-FT (Wortsman et al., 2022) 70.1 31.9 25.3 56.3 79.8 29.9 74.9 45.6 56.8 52.3
   iCaRL (Rebuffi et al., 2017b) 71.7 35.0 43.0 63.4 86.9 43.9 87.8 63.7 60.0 61.7
   ZSCL (Zheng et al., 2023) 73.3 32.6 36.8 62.1 83.8 42.1 83.6 56.5 60.2 59.0
@@
   LwF (Li & Hoiem, 2017) 24.7 79.7 38.3 36.9 63.9 81.0 36.5 71.9 42.7 56.7 53.2
   WiSE-FT (Wortsman et al., 2022) 27.1 76.5 40.9 31.3 68.7 81.6 31.4 74.7 51.7 58.4 54.2
   iCaRL (Rebuffi et al., 2017b) 25.4 72.1 37.5 51.6 65.1 87.1 59.1 88.0 63.7 60.1 61.0
   ZSCL (Zheng et al., 2023) 36.0 75.0 40.7 40.5 71.0 85.3 46.3 83.3 60.7 61.5 60.0
@@
   LwF (Li & Hoiem, 2017) 25.5 72.1 38.9 55.4 65.5 87.3 81.9 88.6 63.6 61.5 64.0
   WiSE-FT (Wortsman et al., 2022) 21.8 76.8 42.9 20.8 77.5 84.9 30.7 76.6 75.8 72.5 58.0
   iCaRL (Rebuffi et al., 2017b) 25.5 72.1 38.9 55.4 65.5 87.3 81.9 88.6 63.6 61.5 64.0
   ZSCL (Zheng et al., 2023) 33.1 75.3 43.5 35.2 74.6 87.4 50.4 84.2 77.3 73.4 63.4
@@
4 Experiments
4.1 Experimental Setting
Datasets. We evaluate our method on two widely used continual learning settings for vision-language models: MTIL (Zheng et al., 2023; Yu et al., 2024) and X-TAIL (Xu et al., 2024). A wide range of fine-grained datasets is used, where each dataset is treated as an individual task to be incrementally learned. We leave detailed experimental settings in Appendix B.
Evaluation Metrics. We use the same metrics as prior work (Zheng et al., 2023; Yu et al., 2024; Xu et al., 2024): “Transfer” for zero-shot performance on unseen data, “Last” for retention of earlier tasks, and “Average” for mean accuracy across all datasets of all learning tasks.
Implementation Details. We follow the setups in (Zheng et al., 2023; Yu et al., 2024; Xu et al., 2024) using CLIP with a ViT-B/16 backbone (Radford et al., 2021). By default, CoDyRA is applied to all pre-trained weight matrices in the vision and text encoders with an initial rank of 16. Each task is trained for 500 iterations with AdamW (Loshchilov & Hutter, 2017), using dense training ratios of 0.5 for X-TAIL and 0.7 for MTIL. $\kappa_{\text{max}}$ is set as 0.005.
4.2 Experimental Results
Cross-domain task-agnostic incremental learning. We report X-TAIL results in Table 1. Unlike prior methods (Yu et al., 2024; Xu et al., 2024) that rely on zero-shot predictions to preserve pre-trained capabilities, our approach continually updates the model with minimal forgetting. As those methods are bounded by zero-shot performance, ours surpasses this limit, achieving higher transfer accuracy. While continual updates may slightly affect some unseen predictions, the overall effect is positive: the model adapts to new tasks while modestly improving generalization.
[…]
Method Average Transfer Accuracy on Unseen Data
CIFAR100 Places365 ImageNet-1k Average
Zero-Shot (Radford et al., 2021) 68.24 33.77 66.72 56.24
MoE-Adapter (Yu et al., 2024) 68.24 33.77 66.72 56.24
RAIL (Xu et al., 2024) 68.24 33.77 66.72 56.24
CoDyRA 68.95 36.52 68.52 58.00
Table 3: Comparisons on 5-shot MTIL setting.
Method
Aircraft Caltech101 CIFAR100 DTD EuroSAT Flowers Food MNIST OxfordPet Cars SUN397 Average